# Modified Lepage-type test statistics for the weak null hypothesis

### v1

@@ 2. $\bullet$
Using a permutation-based approach, we calculated the exact critical values for the $C^{*}$ and $C_{P}^{*}$ test statistics. Our results indicated that both statistics have identical critical values across various sample size combinations. Specifically, the critical values are found at 5% upper-tailed to be 1.697056 for $m=n=5$, 1.669331 for $m=n=10$, 1.674574 for $m=20,n=10$, and 1.62453 for $m=n=20$. The observed similarity in critical values suggests that the asymptotic normality of the $C_{P}^{*}$ test statistic under the null hypothesis is consistent with that of the $C^{*}$ test statistic.
3.4 Foundation in Slutsky’s theorem
Slutsky’s theorem is a foundational principle in probability theory that is instrumental in establishing the asymptotic normality of a statistic, particularly when it incorporates a consistent estimator. The theorem states that if a sequence of random variables, $A_{n}$, converges in distribution, $D$, and another sequence of random variables, $B_{n}$, converges in probability to a constant, $a$, then the product of these sequences, $A_{n}B_{n}$, converges in distribution to $aD$ (see e.g., Casella and Berger 2002, p. 239–240). This can be expressed as
$\text{If}~~A_{n}\overset{d}{\longrightarrow}D~~\text{and}~~B_{n}\overset{p}{\longrightarrow}a,~~\text{then}~~A_{n}B_{n}\overset{d}{\longrightarrow}aD.$
We can apply this theorem to demonstrate the asymptotic normality of the test statistic, $C_{P}^{*}$. This statistic can be expressed as the product of two components in the following form
$C_{P}^{*}=\frac{C-E_{0}(C)}{\sqrt{\widehat{Var}(C)}}=\frac{C-E_{0}(C)}{\sqrt{Var_{0}(C)}}\times\sqrt{\frac{Var_{0}(C)}{\widehat{Var}(C)}}.$ (9)
The first component of Eq. (9), $\frac{C-E_{0}(C)}{\sqrt{Var_{0}(C)}}$, converges in distribution to a standard normal distribution, whereas the second component, $\sqrt{\frac{Var_{0}(C)}{\widehat{Var}(C)}}$, converges in probability to 1. According to Slutsky’s theorem, their product, the test statistic $C_{P}^{*}$ is asymptotically normally distributed under a weak $H_{0}$
$C_{P}^{*}=\frac{C-E_{0}(C)}{\sqrt{\widehat{Var}(C)}}\overset{d}{\longrightarrow}\mathcal{N}(0,1).$
@@ 4.1 Hypotheses
Let $X=(X_{1},X_{2},\ldots,X_{m})$ and $Y=(Y_{1},Y_{2},\ldots,Y_{n})$ denote two i.i.d. random samples of sizes $m$ and $n$, respectively. Let $\mathbb{F}$ and $\mathbb{G}$ represent the continuous distribution functions associated with the samples $X$ and $Y$, respectively. The null hypothesis under consideration posits that both samples are drawn from the same population as
$H_{0}:\mathbb{F}(\theta)=\mathbb{G}(\theta),~~~\forall~\theta\in(-\infty,\infty).$
Our aim is to assess the null hypothesis against the location-scale alternative
$H_{1}:\mathbb{F}(\theta)=\mathbb{G}\left(\frac{\theta+\Delta}{\tau}\right),$
where
$\Delta\neq 0,~~~\tau\neq 1,~~~\Delta\in(-\infty,\infty),~~~\tau\in(0,\infty).$
In hypothesis $H_{1}$, $\Delta$ represents the location difference, while $\tau$ represents the scale difference between the distributions $\mathbb{F}$ and $\mathbb{G}$.
@@ 4.2 The Lepage-type test statistics
The classical Lepage statistic, introduced by Lepage (1971), combines the Wilcoxon-Mann-Whitney UU–test for location difference and the Ansari-Bradley CC–test for scale variation. The $U$ statistic, defined in Eq. (1), is applicable to the location model $\mathbb{F}(\theta)=\mathbb{G}(\theta+\Delta)$, as formalized by the following hypotheses
$H_{0}:\Delta=0~~~~~~\text{versus}~~~~~~H_{1}:\Delta<0,~\Delta>0~~\text{or}~~\Delta\neq 0.$
The CC–test, within the framework of a location-scale model, relates to the scale model $\mathbb{F}(\theta)=\mathbb{G}(\theta/\tau)$ by imposing the restriction that the location parameter $\Delta$ equals zero. This restriction gives the subsequent hypotheses
$H_{0}:\tau=1~~~~~~\text{versus}~~~~~~H_{1}:\tau<1,~\tau>1~~\text{or}~~\tau\neq 1.$
Then Lepage statistic is defined as
$L_{0}=\frac{\big(U-E_{0}(U)\big)^{2}}{Var_{0}(U)}+\frac{\big(C-E_{0}(C)\big)^{2}}{Var_{0}(C)}.$
Lepage demonstrated that the statistics $U$ and $C$ are linearly uncorrelated inside the $L_{0}$ statistic framework for all $m$ and $n$; hence, this test under $H_{0}$ asymptotically conforms to a $\chi^{2}$ distribution with 2 degrees of freedom (Lepage, 1971).
This research presents a methodology for constructing five Lepage-type test statistics to assess a weak null hypothesis. These statistics are constructed using the two estimators for $Var_{0}(U)$ detailed in Section 2 and the single estimator for $Var_{0}(C)$ described in Section 3. These estimators are formulated within the framework of the Lepage test statistic as follows
$L_{1}=\frac{\big(U-E_{0}(U)\big)^{2}}{\widehat{Var}_{FP}(U)}+\frac{\big(C-E_{0}(C)\big)^{2}}{Var_{0}(C)}$
$L_{2}=\frac{\big(U-E_{0}(U)\big)^{2}}{\widehat{Var}_{FH}(U)}+\frac{\big(C-E_{0}(C)\big)^{2}}{Var_{0}(C)}$
$L_{3}=\frac{\big(U-E_{0}(U)\big)^{2}}{Var_{0}(U)}+\frac{\big(C-E_{0}(C)\big)^{2}}{\widehat{Var}(C)}$
$L_{4}=\frac{\big(U-E_{0}(U)\big)^{2}}{\widehat{Var}_{FP}(U)}+\frac{\big(C-E_{0}(C)\big)^{2}}{\widehat{Var}(C)}$
and
$L_{5}=\frac{\big(U-E_{0}(U)\big)^{2}}{\widehat{Var}_{FH}(U)}+\frac{\big(C-E_{0}(C)\big)^{2}}{\widehat{Var}(C)}.$
Under the null hypothesis ($H_{0}$), the test statistic $L_{i}$ (for all $i=1,2,\cdots,5$) is assumed to follow an asymptotic chi-squared distribution with two degrees of freedom. This is because the constituent statistics, $U$ and $C$, are independent of each other and each of them asymptotically follows a standard normal distribution, as discussed in Sections 2 and 3.
@@ 4.3 Critical values for small sample sizes
The Lepage-type test statistics, when used with a normal approximation, are expected to maintain the nominal Type I error rate under the null hypothesis, provided that the sample sizes are sufficiently large. However, for small sample sizes, a permutation-based methodology offers a more robust approach for controlling the Type I error rate. To construct the permutation distribution, an ideal approach would be to enumerate all possible ways of reassigning group labels among the $N$ total samples and calculate the test statistic for each permutation. Due to the computational infeasibility of exhaustive enumeration for even moderately sized samples, a Monte Carlo approach is typically employed. This method involves generating a large number of random permutations to approximate the null distribution.
In this study, we utilized the R package NSM3, which is specifically designed to provide exact and Monte Carlo permutation-based methods for various nonparametric statistical tests, including the Lepage test. The function $cLepage(\alpha,m,n)$” is used to compute the critical values. For all modified Lepage test statistics, we conducted a simulation study under small-sample scenario, with 100,000 simulations. To ensure the consistency of the results, we varied the underlying probability distributions under the null hypothesis, including the normal, uniform, logistic, and Laplace distributions. Across all distributions, the resulting critical values remained consistent. These values are presented in Table 2.
@@
10 10 5.7436 6.5368 6.5719 6.5545 7.5905 7.5905
5 Size and power comparison
A Monte Carlo simulation study is conducted in R (version 4.5.1) to assess the performance of Lepage-type test statistics. Power levels are estimated based on 10,000 simulations, with the significance level for rejecting the null hypothesis ($H_{0}$) set at $0.05$. The power performances of the proposed $L_{i}$ (for all $i=1,2,\cdots,5$) test statistics are compared to that of the classical $L_{0}$ test under simultaneous location-scale shift. For the sake of brevity, the significance level and estimated power of $L_{0}$ and $L_{i}$ (for all $i=1,2,\cdots,5$) are examined for two sample size combinations, $(m,n)$: $(10,10)$ and $(40,30)$. The hypotheses under consideration are $H_{0}:\mathbb{F}(\theta)=\mathbb{G}(\theta)$ versus $H_{1}:\mathbb{F}(\theta)=\mathbb{G}(\frac{\theta-\Delta}{\tau})$, where $\Delta\neq 0$ and $\tau\neq 1$. To evaluate the performance of the competing Lepage-type test statistics, $\mathbb{F}$ and $\mathbb{G}$ are used to represent the following distributions.
@@ 1. $\bullet$
    $\mathbb{F}\sim\mathcal{E}(0.5)$ versus $\mathbb{G}\sim\mathcal{E}(\theta)$: The exponential distribution is a continuous probability distribution characterized by its right-skewed nature. It is frequently employed to model the duration of time until a specific event occurs, such as the time until a mechanical component experiences failure, the inter-arrival time of customers at a service point, or the decay time of a radioactive particle.
@@ 2. $\bullet$
    $\mathbb{F}\sim\chi^{2}_{(1)}$ versus $\mathbb{G}\sim\chi^{2}_{(\theta)}$: The chi-square distribution is a continuous, asymmetric probability distribution characterized by a rightward skew. It is a fundamental tool in inferential statistics, particularly for applications in hypothesis testing.
@@ 3. $\bullet$
    $\mathbb{F}\sim\mathcal{G}(1,1)$ versus $\mathbb{G}\sim\mathcal{G}(1,\theta)$: The Gamma distribution is a versatile, continuous probability distribution characterized by a rightward skew. It is widely employed to model waiting times and other non-negative random variables, finding extensive application in diverse fields, such as reliability engineering, queueing theory, insurance and hydrology, etc.
@@ 4. $\bullet$
    $\mathbb{F}\sim\mathcal{B}(2,2)$ versus $\mathbb{G}\sim\mathcal{B}(2,\theta)$: The beta distribution is a family of continuous probability distributions defined on the interval $[0,1]$. It is primarily utilized to model the behavior of random variables representing proportions, percentages, or probabilities. Unlike unimodal distributions like the normal or Laplace distributions, the beta distribution exhibits remarkable flexibility in its shape. Its form can vary significantly, ranging from a straight line to U-shaped, J-shaped, or even bell-shaped, depending on the specific values of its two parameters. This adaptability makes it a valuable tool in various fields. The beta distribution is well-suited for several applications, including Bayesian inference, project management and reliability engineering.
@@ 5. $\bullet$
    $\mathbb{F}\sim\mathcal{U}(0,1)$ versus $\mathbb{G}\sim\mathcal{U}(0,\theta)$: A continuous uniform distribution is a probability distribution where every outcome within a defined interval has an equal probability of occurrence. This core property makes it a valuable baseline model in statistical analysis and simulations. The uniform distribution is widely used in situations where there is no a priori reason to believe that any single outcome is more likely than another. It is widely employed for statistical analysis, simulations, and the modeling of real-world phenomena.
Simulation results for the exponential distribution are presented in Table 3, examining both small $(m=n=10)$ and large $(m=40,n=30)$ sample sizes. The first column, which corresponds to the case of identical distributions, shows the simulated Type I error rate. A size correct test must have a Type I error rate close to the nominal significance level of $0.05$. For large samples, the $L_{0}$ test appears slightly conservative, while the proposed tests are slightly liberal. This is likely due to the use of asymptotic critical values rather than permutation-based ones. All other columns demonstrate that the proposed tests have significantly higher power than $L_{0}$, with the exception of the $L_{3}$ test. These results suggest that the proposed Lepage-type tests are a suitable alternative for analyzing exponentially distributed data.
Simulation results for the chi-square distribution are displayed in Table 4. For small samples, the Type I error rates are close to the nominal level. However, for large samples, a slightly liberal behavior is observed in the $L_{3}$, $L_{4}$ and $L_{5}$ test statistics. The power comparison shows that the proposed tests consistently maintain higher power across all findings for both small and large sample sizes. As observed previously with the exponential distribution, the proposed $L_{3}$ test statistic demonstrates the lowest power.
An analysis of the Gamma distribution is presented in Table 5. The simulation study indicates that the proposed statistics consistently outperform the classic $L_{0}$ statistic for both small and large sample comparisons, with the exception of $L_{3}$. For Type I error, all tests maintain the rate near $0.05$, though $L_{4}$ and $L_{5}$ exhibit a slightly liberal behavior in the large sample comparison. In summary, the proposed test statistics generally demonstrate greater efficiency than the $L_{0}$ statistic for estimating parameters of the Gamma distribution under both small and large sample conditions.
@@
Table 3: Simulated actual Type I error rate and power of $\mathcal{E}(0.5)$ versus $\mathcal{E}(\theta)$.
@@
Table 4: Simulated actual Type I error rate and power of $\chi^{2}_{(1)}$ versus $\chi^{2}_{(\theta)}$.
$\theta$
$m$ $n$ Test 1.0 1.5 2.0 2.5 3.0
10 10 $L_{0}$ 0.0534 0.1051 0.2626 0.4508 0.6387
$L_{1}$ 0.0492 0.1279 0.3203 0.5284 0.7110
$L_{2}$ 0.0496 0.1373 0.3124 0.5320 0.7165
$L_{3}$ 0.0503 0.1022 0.2135 0.3814 0.5606
$L_{4}$ 0.0536 0.1188 0.2726 0.4802 0.6525
$L_{5}$ 0.0511 0.1146 0.2831 0.4747 0.6642
40 30 $L_{0}$ 0.0480 0.3230 0.7984 0.9738 0.9978
$L_{1}$ 0.0535 0.3582 0.8320 0.9811 0.9986
$L_{2}$ 0.0521 0.3492 0.8265 0.9806 0.9987
$L_{3}$ 0.0584 0.3348 0.7925 0.9772 0.9985
$L_{4}$ 0.0604 0.3709 0.8323 0.9815 0.9987
$L_{5}$ 0.0583 0.3659 0.8309 0.9790 0.9984
@@
Table 5: Simulated actual Type I error rate and power of $\mathcal{G}(1,1)$ versus $\mathcal{G}(1,\theta)$.
$\theta$
$m$ $n$ Test 1.0 2.0 3.0 4.0 5.0
10 10 $L_{0}$ 0.0496 0.1882 0.4070 0.6077 0.7246
$L_{1}$ 0.0490 0.2236 0.4787 0.6627 0.7916
$L_{2}$ 0.0475 0.2220 0.4770 0.6621 0.7812
$L_{3}$ 0.0504 0.1638 0.3523 0.5333 0.6714
$L_{4}$ 0.0533 0.1951 0.4282 0.6200 0.7492
$L_{5}$ 0.0478 0.1962 0.4241 0.6216 0.7501
40 30 $L_{0}$ 0.0496 0.6163 0.9621 0.9971 0.9997
$L_{1}$ 0.0500 0.6651 0.9712 0.9985 1.0000
$L_{2}$ 0.0494 0.6504 0.9677 0.9989 1.0000
$L_{3}$ 0.0536 0.6255 0.9633 0.9975 0.9999
$L_{4}$ 0.0607 0.6713 0.9716 0.9982 0.9999
$L_{5}$ 0.0612 0.6735 0.9731 0.9979 0.9999
In a simulation study of the beta distribution, the proposed test statistics clearly outperform the traditional $L_{0}$ statistic in detecting simultaneous shifts under both small and large sample comparisons, as shown in Table 6. The $L_{4}$ and $L_{5}$ tests are slightly liberal in the large sample comparison. The proposed tests show superior power compared to the classical Lepage test. For this distribution, $L_{3}$ has the lowest power in the small sample investigation but performs similarly to $L_{0}$ in the large sample case. These findings suggest that the proposed tests offer a robust alternative to $L_{0}$ for analyzing parameters of the beta distribution.
For the uniform distribution, all competing tests show similar performance in both sample scenarios, see Table 7. Consistent with previous findings, the proposed $L_{3}$, $L_{4}$, and $L_{5}$ tests are slightly liberal for large sample investigation.
@@
Table 6: Simulated actual Type I error rate and power of $\mathcal{B}(2,2)$ versus $\mathcal{B}(2,\theta)$.
$\theta$
$m$ $n$ Test 2.0 3.0 4.0 5.0 6.0
10 10 $L_{0}$ 0.0495 0.1221 0.2927 0.4647 0.6441
$L_{1}$ 0.0512 0.1414 0.3333 0.5486 0.7104
$L_{2}$ 0.0504 0.1391 0.3413 0.5482 0.7093
$L_{3}$ 0.0487 0.1033 0.2407 0.3971 0.5708
$L_{4}$ 0.0518 0.1252 0.3074 0.4871 0.6692
$L_{5}$ 0.0503 0.1204 0.3059 0.4910 0.6614
40 30 $L_{0}$ 0.0509 0.3688 0.8383 0.9843 0.9989
$L_{1}$ 0.0512 0.4127 0.8698 0.9874 0.9981
$L_{2}$ 0.0511 0.4020 0.8659 0.9852 0.9990
$L_{3}$ 0.0525 0.3737 0.8420 0.9837 0.9988
$L_{4}$ 0.0633 0.4142 0.8763 0.9855 0.9992
$L_{5}$ 0.0577 0.4188 0.8683 0.9856 0.9990

### v3

@@ 4.1 Hypotheses
Let $X=(X_{1},\ldots,X_{m})\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathbb{F}$ and $Y=(Y_{1},\ldots,Y_{n})\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathbb{G}$ be two independent samples, where $\mathbb{F}$ and $\mathbb{G}$ represent continuous distribution functions. We consider two frameworks for hypothesis testing
Strong null hypothesis:
$H_{0}^{S}:\mathbb{F}(x)=\mathbb{G}(x)\quad\forall x\in(-\infty,\infty),$
against the location-scale alternative
$H_{1}:\mathbb{F}(x)=\mathbb{G}\left(\frac{x-\theta_{1}}{\theta_{2}}\right),\quad\theta_{1}\neq 0,\quad\theta_{2}\neq 1.$
Weak null hypothesis:
$H_{0}^{W}:\mu_{1}=\mu_{2}\quad\text{and}\quad\sigma_{1}=\sigma_{2},$
without assuming $\mathbb{F}=\mathbb{G}$, which allows for potential heteroscedasticity under the null. The classical Lepage test is designed for the strong null hypothesis $H_{0}^{S}$, while the proposed robustified tests target the more realistic weak null hypothesis $H_{0}^{W}$.
@@ 4.2 The Lepage-type test statistics
The classical Lepage statistic, proposed by Lepage (1971), combines the $U$-test statistic for location difference and the $C$-test statistic for scale variation
$L_{0}=\frac{(U-E_{0}(U))^{2}}{Var_{0}(U)}+\frac{(C-E_{0}(C))^{2}}{Var_{0}(C)}.$ (10)
Under the strong null hypothesis ($H_{0}^{S}$), Lepage established the asymptotic independence of $U$ and $C$, ensuring that the Lepage statistic $L_{0}$ follows an asymptotic $\chi^{2}_{2}$ distribution. In this research, we propose five new test statistics that incorporate robust variance estimators. The formulations of these statistics are as follows
$L_{1}$ $=\frac{(U-E_{0}(U))^{2}}{\widehat{Var}_{FP}(U)}+\frac{(C-E_{0}(C))^{2}}{Var_{0}(C)}$
$L_{2}$ $=\frac{(U-E_{0}(U))^{2}}{\widehat{Var}_{FH}(U)}+\frac{(C-E_{0}(C))^{2}}{Var_{0}(C)}$
$L_{3}$ $=\frac{(U-E_{0}(U))^{2}}{Var_{0}(U)}+\frac{(C-E_{0}(C))^{2}}{\widehat{Var}(C)}$
$L_{4}$ $=\frac{(U-E_{0}(U))^{2}}{\widehat{Var}_{FP}(U)}+\frac{(C-E_{0}(C))^{2}}{\widehat{Var}(C)}$
$L_{5}$ $=\frac{(U-E_{0}(U))^{2}}{\widehat{Var}_{FH}(U)}+\frac{(C-E_{0}(C))^{2}}{\widehat{Var}(C)}.$
4.3 Theoretical properties
Proposition 2 (Asymptotic distribution)
Under the weak null hypothesis $H_{0}^{W}$ and assuming continuous distributions with finite moments, each test statistic $L_{i}$ ($i=1,\ldots,5$) converges in distribution to a chi-square distribution with 2 degrees of freedom.
Proof
For each $L_{i}$, we can write
$L_{i}=T_{1i}^{2}+T_{2i}^{2},$
where $T_{1i}=(U-E_{0}(U))/\sqrt{\widehat{Var}_{i}(U)}$ and $T_{2i}=(C-E_{0}(C))/\sqrt{\widehat{Var}_{i}(C)}$ are the standardized components using the appropriate variance estimators. From Sections 2 and 3, each $T_{1i}$ and $T_{2i}$ converges in distribution to standard normal random variables under $H_{0}^{W}$. While the independence of $U$ and $C$ is not guaranteed under the weak null, the quadratic form construction remains valid as long as the components are asymptotically normal. The asymptotic $\chi^{2}_{2}$ distribution follows from the continuous mapping theorem. $\square$
@@ 4.4 Critical values for small sample sizes
To ensure accurate Type I error control for small samples, we obtained critical values via a permutation approach. This was implemented using the cLepage() function from the NSM3 package (Schneider et al., 2025) in R (version 4.5.1) with 100,000 Monte Carlo permutations under the strong null hypothesis. The critical values, detailed in Table 2, were found to be consistent across the normal, uniform, logistic, and Laplace distributions. For larger samples ($m,n>30$), the asymptotic $\chi^{2}_{2}$ distribution provides adequate approximation, though permutation tests remain recommended for precise Type I error control.
Remark 1
The critical values for statistics $L_{1}$ and $L_{2}$ (and analogously, for $L_{4}$ and $L_{5}$) coincide when the sample size $N$ is even. This empirical observation is corroborated by the simulation results presented in Section 5, which demonstrate that these statistic pairs frequently yield nearly identical outcomes. This phenomenon arises because the Fligner-Policello and Fong-Huang variance estimators, despite their differing theoretical constructions, often converge to similar numerical estimates in finite samples. Consequently, for practical applications, the choice between these estimators may be guided by computational considerations or the specific context of the problem, as their performance is largely comparable.
@@
10 10 5.7436 6.5719 6.5719 6.5545 7.5905 7.5905
5 Simulation study: size and power comparison
A comprehensive Monte Carlo simulation study was conducted in R to evaluate the Type I error control and power of the proposed tests. Empirical sizes and powers were estimated based on 10,000 replications at a nominal significance level of $\alpha=0.05$. The performance of the proposed tests $L_{1}$ to $L_{5}$ was compared against the classical $L_{0}$ test under various location-scale shifts. For the sake of brevity, results are presented for two sample size combinations: a small, balanced design $(m,n)=(10,10)$ and a larger, unbalanced design $(m,n)=(40,30)$. To evaluate the performance of the competing Lepage-type test statistics, $\mathbb{F}$ and $\mathbb{G}$ were specified as the following right-skewed distributions:
@@ 1. $\bullet$
    Exponential: $\mathbb{F}\sim Exp(0.5)$ versus $\mathbb{G}\sim Exp(\theta)$. The exponential distribution is a continuous probability distribution characterized by its right-skewed nature. A change in the scale parameter $\theta$ induces a proportional shift in both the location and scale of the distribution.
@@ 2. $\bullet$
    Gamma: $\mathbb{F}\sim G(2,2)$ versus $\mathbb{G}\sim G(\theta_{1},\theta_{2})$. The Gamma distribution is a versatile, continuous probability distribution characterized by a rightward skew.
@@ 3. $\bullet$
    Chi-square: $\mathbb{F}\sim\chi^{2}_{2}$ versus $\mathbb{G}\sim\theta_{2}\cdot\chi^{2}_{2}+\theta_{1}$. The $\chi^{2}$ distribution is a continuous, asymmetric probability distribution characterized by a rightward skew. The parameters $\theta_{1}$ and $\theta_{2}$ induce location and scale shifts, respectively.
@@ 4. $\bullet$
    Lognormal: $\mathbb{F}\sim LN(0,2)$ versus $\mathbb{G}\sim LN(\theta_{1},\theta_{2})$. The lognormal distribution is a continuous probability distribution characterized by a rightward skew. The parameters $\theta_{1}$ and $\theta_{2}$ induce location and scale shifts, respectively.
@@ 5. $\bullet$
    Weibull: $\mathbb{F}\sim W(2,1)$ versus $\mathbb{G}\sim W(\theta_{1},\theta_{2})$. The Weibull distribution is a versatile continuous distribution characterized by its flexibility in modeling various failure rates.
5.1 Type I error control
The first column of Tables 3 to 7, corresponding to the case of identical distributions ($\theta=0.5$ for exponential; $\theta_{1}=2,\theta_{2}=2$ for gamma; etc.), shows the simulated Type I error rates. For the small sample size $(m,n)=(10,10)$, all competing tests, including the classical $L_{0}$ and the proposed $L_{1}$$L_{5}$, generally maintain their empirical size within or very close to the acceptable range around the nominal $0.05$ level. For instance, in the Exponential distribution (Table 3), the empirical sizes range from 0.0502 to 0.0526.
[…]
5.2 Power performance
The subsequent columns in Tables 3 to 7 demonstrate the empirical power of the tests under various alternatives. The results reveal a consistent and clear pattern across all five right-skewed distributions.
1. Superior power of proposed tests: The proposed tests $L_{1}$, $L_{2}$, $L_{4}$, and $L_{5}$ almost universally achieve higher power than the classical $L_{0}$ test. For example, in the exponential distribution with $(m,n)=(10,10)$ and $\theta=1.5$ (Table 3), the power of $L_{0}$ is 0.4216, while $L_{1}$ and $L_{2}$ achieve a power of 0.4799—a substantial increase. This superiority is maintained in larger samples; with $(m,n)=(40,30)$ and $\theta=1.0$, $L_{1}$ and $L_{2}$ (power $\approx$0.66) outperform $L_{0}$ (power 0.6107).
2. The critical role of both components: The test $L_{3}$, which modifies only the scale ($C$) component, generally shows the lowest power among the proposed tests and often performs similarly to or even worse than $L_{0}$, particularly for location-dominated shifts. This indicates that improving the location ($U$) component is crucial for overall power gain. Conversely, the tests that modify both components ($L_{4}$ and $L_{5}$) consistently most powerful than the traditional $L_{0}$, demonstrating that the full benefit is realized when both the location and scale tests are robustified.
3. Performance hierarchy: Based on the aggregate results, a discernible performance hierarchy emerges: $L_{5}\approx L_{4}\gtrsim L_{2}\approx L_{1}>L_{0}>L_{3}$. The tests incorporating the Fong-Huang (FH) estimator for the $U$ statistic ($L_{2}$ and $L_{5}$) often show a slight, consistent advantage over those using the Fligner-Policello (FP) estimator ($L_{1}$ and $L_{4}$). For instance, in the lognormal distribution with $(m,n)=(40,30)$ and $(\theta_{1}=1,\theta_{2}=1.5)$ (Table 6), $L_{5}$ and $L_{4}$ have marginally higher power than the other tests (power $\approx$0.67).
4. Robustness to distributional form: The power advantage of the proposed tests is not an artifact of a single distribution. It is robustly observed across the exponential, gamma, chi-square, lognormal, and Weibull distributions. This confirms that the proposed enhancements are broadly effective for the analysis of right-skewed data, which is the primary focus of this work.
In summary, the simulation study confirms that the proposed Lepage-type tests, particularly $L_{2}$ and $L_{5}$, offer a superior alternative to the classical $L_{0}$ test. They provide significantly higher power for detecting location-scale shifts in right-skewed data while maintaining reasonable control over the Type I error rate. The slight liberal tendency of $L_{3}$, $L_{4}$, and $L_{5}$ in larger samples is a minor trade-off for their substantial power gains, and this can be mitigated in practice by using a permutation approach for critical value determination.
@@ 5.2 Power performance
Table 3: Simulated actual Type I error rate and power of $Exp(0.5)$ versus $Exp(\theta)$.
@@
Table 4: Simulated actual Type I error rate and power of $G(2,2)$ versus $G(\theta_{1},\theta_{2})$.
$\theta_{2}$
$m$ $n$ $\theta_{1}$ Test 2.0 1.5 1.0 0.5
10 10 2.0 $L_{0}$ 0.0512 0.0994 0.3596 0.9159
$L_{1}$ 0.0487 0.1138 0.4202 0.9456
$L_{2}$ 0.0487 0.1138 0.4202 0.9456
$L_{3}$ 0.0547 0.0884 0.3016 0.8797
$L_{4}$ 0.0521 0.1029 0.3800 0.9276
$L_{5}$ 0.0521 0.1029 0.3800 0.9276
3.0 $L_{0}$ 0.2116 0.4899 0.8782 0.9990
$L_{1}$ 0.2604 0.5753 0.9209 0.9995
$L_{2}$ 0.2604 0.5753 0.9209 0.9995
$L_{3}$ 0.1737 0.4149 0.8297 0.9977
$L_{4}$ 0.2274 0.5236 0.8975 0.9994
$L_{5}$ 0.2274 0.5236 0.8975 0.9994
40 30 2.0 $L_{0}$ 0.0495 0.2701 0.9283 1.0000
$L_{1}$ 0.0543 0.2936 0.9374 1.0000
$L_{2}$ 0.0544 0.2938 0.9375 1.0000
$L_{3}$ 0.0568 0.2715 0.9278 1.0000
$L_{4}$ 0.0583 0.2960 0.9373 1.0000
$L_{5}$ 0.0582 0.2963 0.9375 1.0000
3.0 $L_{0}$ 0.6837 0.9834 1.0000 1.0000
$L_{1}$ 0.7243 0.9872 1.0000 1.0000
$L_{2}$ 0.7244 0.9874 1.0000 1.0000
$L_{3}$ 0.6869 0.9834 1.0000 1.0000
$L_{4}$ 0.7271 0.9875 1.0000 1.0000
$L_{5}$ 0.7273 0.9876 1.0000 1.0000
@@
Table 5: Simulated actual Type I error rate and power of $\chi^{2}_{2}$ versus $\theta_{2}*\chi^{2}_{2}+\theta_{1}$.
$\theta_{2}$
$m$ $n$ $\theta_{1}$ Test 1.0 1.5 2.0 2.5
10 10 0.0 $L_{0}$ 0.0475 0.0959 0.1939 0.2977
$L_{1}$ 0.0479 0.1031 0.2274 0.3459
$L_{2}$ 0.0479 0.1031 0.2274 0.3459
$L_{3}$ 0.0526 0.0855 0.1652 0.2517
$L_{4}$ 0.0488 0.0963 0.2012 0.3100
$L_{5}$ 0.0488 0.0963 0.2012 0.3100
1.0 $L_{0}$ 0.3672 0.6701 0.8779 0.9599
$L_{1}$ 0.3751 0.7197 0.9141 0.9756
$L_{2}$ 0.3751 0.7197 0.9141 0.9756
$L_{3}$ 0.3358 0.5971 0.8248 0.9362
$L_{4}$ 0.3556 0.6752 0.8849 0.9661
$L_{5}$ 0.3556 0.6752 0.8849 0.9661
40 30 0.0 $L_{0}$ 0.0469 0.2534 0.6339 0.8683
$L_{1}$ 0.0521 0.2709 0.6560 0.8801
$L_{2}$ 0.0521 0.2706 0.6553 0.8794
$L_{3}$ 0.0550 0.2558 0.6331 0.8670
$L_{4}$ 0.0594 0.2741 0.6542 0.8780
$L_{5}$ 0.0594 0.2740 0.6534 0.8778
1.0 $L_{0}$ 0.9339 0.9996 1.0000 1.0000
$L_{1}$ 0.9431 0.9997 1.0000 1.0000
$L_{2}$ 0.9434 0.9997 1.0000 1.0000
$L_{3}$ 0.9475 0.9995 1.0000 1.0000
$L_{4}$ 0.9545 0.9997 1.0000 1.0000
$L_{5}$ 0.9546 0.9998 1.0000 1.0000
@@
Table 6: Simulated actual Type I error rate and power of $LN(0,2)$ versus $LN(\theta_{1},\theta_{2})$.
$\theta_{2}$
$m$ $n$ $\theta_{1}$ Test 2.0 1.5 1.0 0.5
10 10 0.0 $L_{0}$ 0.0496 0.0825 0.2457 0.6972
$L_{1}$ 0.0484 0.0749 0.1919 0.6025
$L_{2}$ 0.0484 0.0749 0.1919 0.6025
$L_{3}$ 0.0506 0.1010 0.3152 0.8031
$L_{4}$ 0.0482 0.0889 0.2677 0.7517
$L_{5}$ 0.0482 0.0889 0.2677 0.7517
1.0 $L_{0}$ 0.1284 0.1900 0.3773 0.7953
$L_{1}$ 0.1551 0.2094 0.3503 0.7252
$L_{2}$ 0.1551 0.2094 0.3503 0.7252
$L_{3}$ 0.1078 0.1729 0.3875 0.8424
$L_{4}$ 0.1341 0.1953 0.3748 0.8049
$L_{5}$ 0.1341 0.1953 0.3748 0.8049
40 30 0.0 $L_{0}$ 0.0489 0.1715 0.7606 0.9991
$L_{1}$ 0.0529 0.1781 0.7630 0.9991
$L_{2}$ 0.0531 0.1783 0.7632 0.9992
$L_{3}$ 0.0553 0.2204 0.8479 0.9997
$L_{4}$ 0.0594 0.2272 0.8494 0.9997
$L_{5}$ 0.0592 0.2274 0.8495 0.9997
1.0 $L_{0}$ 0.3952 0.6098 0.9470 1.0000
$L_{1}$ 0.4292 0.6478 0.9523 1.0000
$L_{2}$ 0.4294 0.6484 0.9525 1.0000
$L_{3}$ 0.3983 0.6311 0.9648 1.0000
$L_{4}$ 0.4321 0.6673 0.9694 1.0000
$L_{5}$ 0.4322 0.6674 0.9695 1.0000